OpenStudy (anonymous):

Find an equation in standard form of the parabola through the points. (1, -4) (2, -3) (3, -4)

OpenStudy (alyssa_xo):

can you pinpoint the vertex?

OpenStudy (nayef):

\[y=ax ^{2}+bx+c << parabola standard form \] Then solve the equations (They are coming from substituting the points in the standard form) \[-4=a+b+c, -3=4a+2b+c, -4=8a+3b+c\].

OpenStudy (nayef):

If you try it , Then I will check your answers :)

OpenStudy (anonymous):

Nayef way is Robotic AND Alyssa's way is "intuitive" and smart. I go with Alyssa. If you plot the points then you can trace the graph of the parabola

OpenStudy (anonymous):

you can tell easily its vertex and then write its vertex form